# RAGGED: Towards Informed Design of Retrieval Augmented Generation Systems

### v1

@@ 6.1 RAGGED Analysis Guidelines
Analyze the reader model’s performance using different slices of instances that represent different qualities of retrieved contexts.
The first slice is where the retrieved context includes at least one gold, human-annotated passage. In this scenario, there is sufficient information in the top-k passages to answer the question. The alternative reader context settings to compare with are when there are a) only the gold passages in the top-k retrieved passages and b) no context. Comparing with a) shows how good the reader is at ignoring noise and how close it is to using only signal and no noise. Comparing with b) shows if the reader is prone to performing worse
The second slice of instances is where none of the top-k retrieved contexts include gold passages. In this scenario, all the passages are noise. The alternative reader context setting to compare with is when there is no context.
@@
Models may show varying degrees of sensitivity to the quality of context. Reader models provided with only gold passages often act as an upper bound for the reader models’ performance with top-k passages. Although one might expect reader models provided with no context passages to act as a lower bound for the reader models’ performance with top-k passages, that is not always the case. It depends on the reader’s ability to discern between potentially sufficient internal knowledge and irrelevant context knowledge.
@@
For practitioners, these comparisons would tell us how effective the reader model sifts for signal among noise. If the model performs very closely to how it would with only gold passages, then it is highly effective at using relevant information and ignoring irrelevant ones. Conversely, a significant drop in performance with lower-quality contexts suggests a reliance on high-quality retrieval and potential vulnerability to noise. This analysis can guide the development or selection of more robust models that can maintain performance despite the variability in context quality.
6.2 No-context Generalization from Pre-trained Knowledge
We evaluate how capable models are at answering questions directly from their parameters and without any context by evaluating reader test-time performance without any context. We denote this setup as no-ctx and compare it with the standard setting with top-1 passages from ColBERT.
| Model | NQ | HotpotQA | BioASQ | |||
| no-ctx | top-1 | no-ctx | top-1 | no-ctx | top-1 | |
| Flan-T5 XXL | 16.7 | 33.8 | 27.4 | 43.5 | 18.2 | 40.0 |
| Flan-UL2 | 23.7 | 35.0 | 29.6 | 43.4 | 19.2 | 30.2 |
| LLaMa2 7B | 21.5 | 29.2 | 23.5 | 38.2 | 21.1 | 32.8 |
| LLaMa2 70B | 34.1 | 35.6 | 32.5 | 46.4 | 26.0 | 35.9 |
As shown in Table 1, under the no-context setting, LLaMa 70B outperforms Flan models by a large margin, suggesting they have at least encoded more knowledge from pretraining or are better at generalizing from memorized knowledge. While LLaMa has potentially memorized more relevant information during training, this knowledge may be more likely to cause conflicts with passages provided in context at test time. Next, we examine two particular slices of examples and study how reader models resolve conflict with irrelevant passages when gold (human-annotated) passages are present (§6.3) or absent (§6.4).
6.3 With Ground-Truth Passages
We study how readers handle context with both positive (gold) passages and negative (non-gold) passages. Specifically, for each given , we first find the slice of examples where the retriever returns at least one gold passage in top- candidates. With this slice of examples, we compare the results when providing (1) top- retrieved passages, (2) only the gold passage(s) within the top-k passages, and (3) no passages at all.
Figure 4 shows that LLaMa models are more severely affected by noisy context, as their scores decrease more sharply as increases. Notably, around and , LLaMa 7B and 70B models’ performance deteriorates to a level below that of their no-context baseline. In comparison, Flan models consistently outperform their no-context counterparts by a large margin.
Special Questions and Domains
We perform similar analyses on HotpotQA and BioASQ datasets. While Flan models exhibit trends akin to their open-domain behaviors, we observe substantial behavioral shifts on LLaMa models (Figure 5). Unlike the severe degradation on NQ questions, LLaMa consistently outperforms the no-context baseline. Additional special-domain contexts seem to induce less conflict with LLaMa’s parametric knowledge than additional contexts do in NQ potentially because they provide new, useful information that the model cannot derive from its pertaining memory alone.
6.4 Response to Negative passages
We now study the setting where no passages are explicitly useful to answer the questions. At each , we select the slice of examples where no gold passages exist among the top- retrieved candidates. We report model results when providing (1) these negative contexts or (2) no contexts at all.
As shown in Figure 6, compared to the no-ctx baseline, LLaMa performance degrades when contextualizing more negative passages (i.e., as increases). On the other hand, Flan models can benefit from negative contexts, consistently outperforming no-ctx baselines across all . Surprisingly, Flan models perform better with negative passages than with contexts at all. One explanation is that the negative passages may still be partially helpful even though they do not contain the exact correct answer since they may still contain some keywords or related topics that provide hints to the reader models. While LLaMa models fail to identify such subtle information, Flan models are more sensitive to related knowledge and capture these cues effectively.
Special Questions
We perform similar analyses on HotpotQA and BioASQ datasets.333For multi-hop questions, we select examples retrieved with all gold passages within the top- passages since all passages are necessary to answer the question. On HotpotQA, Flan models obtain less gain using negative passages and levels off above the no-ctx baseline (Figure 7). In contrast, LLaMa can use negative passages more effectively on special questions, outperforming no-ctx for most ’s on HotpotQA and on BioASQ.
Comparing models of different sizes, smaller models benefit more from negative contexts while larger models suffer. Between the two LLaMa models, while LLaMa 70B significantly regresses beyond its no-context baseline, LLaMa 7B performs comparably with its no-context baseline. Between the two Flan models, while FlanT5 benefits a lot from negative passages and outperforms no-ctx by a large margin, Flan-UL2 becomes sensitive to negative information and deteriorates below its no-ctx baseline across nearly all k values.
Case Study: What Negative Contexts Help?
If a negative context is related but insufficient for answering the question, it can still help improve the reader model’s performance. For instances where FlanT5 answers correctly with the top-5, negative contexts but incorrectly with no contexts, 43% of instances have top-k passages that include at least one passage from the Wikipedia page that the gold passages are from. Even though none of these negative passages include the exact correct answer, they are topically related enough that FlanT5 can bridge the knowledge gap.
Nonetheless, not all negative passages from the gold page are helpful. Among instances where top-5 underperforms no-ctx, 28% include at least one passage from the ground-truth Wikipedia page.
7 Comparing Different Retrievers
@@
We compare retrievers of different quality (§7.2) and study their influence on the reader model’s downstream performance (§7.3).
7.1 Ragged Analysis Guidelines
@@
Retriever performance highly depends on the domain and complexity of the question (number of hops). However, interestingly, good retriever performance does not always guarantee good reader performance. The reader performance also depends on the reader’s sensitivity to retriever quality, with some models showing robustness to suboptimal retrieval.
@@
This aspect of the analysis informs the optimal choice of retriever for your RAG system and how much optimizing for that choice helps the reader’s performance. A model that performs well with various retrievers offers flexibility and might be more versatile in real-world applications. In contrast, a model highly sensitive to retrieval quality might require a state-of-the-art retriever to achieve optimal performance.
7.2 Evaluating Different Retrievers
We evaluate BM25 and ColBERT’s retrieval performance by computing recall@k for all k from 1 to 50 (Table 2). In general, ColBERT performs better than BM25 on questions about Wikipedia knowledge (NQ and HotpotQA). However, ColBERT only offers a negligible advantage over BM25 on the special-domain BioASQ, yet at a much higher computation cost.
Averaged over ranging from 1 to 50, ColBERT offers a 50-point gain over BM25 for NQ, but only offers a point gain over BM25 for BioASQ for paragraph recall. When applying to the out-of-domain questions, ColBERT does not gain many advantages from its neural, contextual embeddings.
| Model | Recall@k | |||||
| 1 | 2 | 5 | 10 | 20 | 50 | |
| NQ | ||||||
| BM25 | 0.7 | 1.1 | 2.5 | 4.1 | 6.0 | 7.5 |
| 10.3 | 16.3 | 27.8 | 36.8 | 47.7 | 53.2 | |
| ColBERT | 12.3 | 18.0 | 25.7 | 32.1 | 38.1 | 41.8 |
| 27.2 | 38.8 | 54.4 | 65.0 | 72.9 | 77.2 | |
| HotpotQA | ||||||
| BM25 | 0.2 | 0.4 | 1.0 | 1.6 | 2.4 | 3.0 |
| 23.3 | 31.2 | 42.7 | 52.1 | 59.1 | 62.8 | |
| ColBERT | 31.1 | 40.1 | 49.9 | 56.2 | 61.9 | 64.9 |
| 34.2 | 44.7 | 56.3 | 63.6 | 69.9 | 73.1 | |
| BioASQ | ||||||
| BM25 | 8.8 | 12.9 | 19.6 | 25.8 | 33.3 | 37.8 |
| 12.4 | 16.4 | 23.9 | 30.6 | 38.7 | 43.6 | |
| ColBERT | 8.8 | 13.5 | 20.7 | 27.1 | 34.3 | 38.6 |
| 14.2 | 18.2 | 25.6 | 32.2 | 39.8 | 44.2 | |
7.3 Generation with Different Retrievers
More importantly, we are interested in the question: how much does the choice of retriever affect downstream performance? We find that the nature of the task (e.g., single-hop vs. multi-hop, domain specificity) can significantly affect the choice between neural and sparse retrievers (Figure 8). While neural retrievers may require more resources, their great advantage in single-hop, Wikipedia-based questions and small advantage in specialized domains may still justify the investment. In contrast, neural retrieves are less beneficial for multi-hop questions, even for the Wikipedia domain.
Open Domain (Wikipedia)
For open-domain single-hop questions, using ColBERT offers substantial improvements over using BM25. At , ColBERT helps Flan models achieve a significant 16–18 points EM improvement and LLaMa2 models a more modest 4–6 point increase. This suggests there is value in using high-quality neural retrieval for open-domain, single-hop questions.
Special Question Type: Multi-Hop
For multi-hop questions, despite the stark retrieval performance gap between BM25 and ColBERT, the impact on reader performance is minimal. At , Flan models improve by 3–4 in F when paired with ColBERT over BM25, while LLaMa2 models improve by 2–3 in F. Both models show marginal gains with ColBERT, suggesting that multi-hop questions challenge reader models more in reasoning than context utilization.
While BM25 consistently gets recall scores 10 points lower than ColBERT, it is surprising that the reader performs comparably to using ColBERT-retrieved passages. This could be explained by how BM25 still has a high Wikipedia page recall, indicating its capability of finding relevant but inexact information that readers may still benefit from.
Special Biomedical Domain
ColBERT and BM25 perform similarly in the biomedical domain, so we might also expect the downstream performances to be similar. However, there is still a difference, albeit a small one, between using ColBERT and BM25. For , Flan models achieve a 2–5 point improvement when paired with ColBERT over being paired with BM25, while LLaMa2 models exhibit a much smaller, point improvement. Although there are only slight differences in retrieval quality, the specialized nature of the domain can amplify the impact on reader performance. As a result, there is still a discernible, though modest, preference for ColBERT over BM25 in accessing quality context.
@@ Appendix A Implementation Details
Specifically, when using FlanT5 and FlanUL2 readers, we use T5Tokenizer to truncate sequences to up to 2 tokens; when using LLaMa2 models, we apply the LlamaTokenizer and truncate sequences by 4 tokens. Subsequently, we incorporate a concise question-and-answer format that segments the query using "Question:" and cues the model’s response with "Answer:", ensuring precise and targeted answers.
For our reader decoding strategy, we used greedy decoding with a beam size of 1 and temperature of 1, selecting the most probabble next word at each step without sampling. The output generation was configured to produce responses with 10 and 256 tokens to examine the effect of varying response lengths on model performance. The experiments were conducted on NVIDIA A6000 GPUs, supported by an environment with 60GB RAM. The average response time was 1.1s per query when processing with a batch size of 50.

### v2

@@ 6 Reader Robustness in the Presence and Absence of Gold Passages
Analyze the reader model’s performance on two slices of instances representing different qualities of retrieved contexts. In §6.1, we demonstrate analyzing the slice where the retrieved passages include at least one gold passage. This represents the scenario where there is sufficient context information to answer the question. In this slice, we compare how the reader performs with a) the top-k passages, b) only the gold passages in the top-k passages (top-gold), and c) no context.
In §6.2, we demonstrate analyzing on the slice of instances, where none of the top-k retrieved passages include any gold passages. This represents the scenario where there is insufficient context information to answer the question. This setting is used to evaluate the model’s ability to ignore noise and fall back onto any helpful pre-trained knowledge.
@@
Reader models provided with only gold passages often serve as an upper bound for their top-k performance. However, reader performance without context does not necessarily represent a lower bound for their top-k performance. This depends on the reader’s ability to distinguish between irrelevant information and helpful pre-trained knowledge.
@@
For practitioners, these analyses can help identify which readers are more robust to noise. A small gap between top-k and top-gold performance means the reader is highly robust to noise and is effective at identifying and using relevant information. Conversely, a significant gap between top-k and top-gold performance suggests the model struggles to discern signal from noise.
6.1 With Gold Passages
For NQ instances where gold passages are in the top-k retrieved passages, GPT-3.5 and Flan models effectively identify and use relevant information, consistently performing better with context than without Figure 4. In contrast, the rest of the models struggle more with noise. For example, Claude Haiku and LLaMa2 models’ top-k performance falls below their no-context performance at and LLaMa3 models at . This highlights how the wrong configuration can result in using RAG being worse than not using RAG even when there is sufficient context information.
For HotpotQA and BioASQ, models generally follow similar trends to NQ regarding which models’ no-context performance serves as a strict lower bound. We provide a detailed discussion of the more interesting models below and include complete plots in Appendix D.
One difference between NQ and HotpotQA performance is that LLaMa2 7B’s top-k performance now dips below the no-context baseline much later at instead of at , and LLaMa2 70B dips below the no-context baseline at instead of at (Figure 5). Similarly, Claude now dips below no-context baseline at instead of . This could suggest that if the problem requires more than one signal passage and all such signals are present, the reader may have more anchor points for the model to rely on before being distracted by the noise in increasing .
Different from for NQ, the gap between top-pos and top-k for BioASQ is much smaller for all models (Figure 10). One potential explanation is that the specialized domain jargon makes it easier for the reader to perform a second-step filtering for signal after the retriever’s first step of filtering. We attribute this difference to the reader’s filtering ability instead of the retriever’s since ColBERT performs strictly worse on BioASQ than on NQ. Another difference is that for Claude Haiku and LLaMa3 70B, even the top-gold dips below the no-context baseline, showing that they struggle much more with special domains. A closer examination reveals that these readers tend to start returning gibberish as the number of contexts increases, even when all the contexts are gold passages.
6.2 Without Gold Passages
For NQ and HotpotQA, almost all models’ top-k performance is worse than their no-context performance (Figure 6, Figure 11). This general trend is not too surprising since the models are instructed to answer based on the context, and these slices of data do not include any gold passages.
What is more surprising is the exception of the Flan models, which benefit even from non-gold contexts and consistently outperform its no-context baseline across all . One potential reason this is possible is that even though the non-gold passages may provide insufficient information, they could still provide relevant keywords and information that act as clues for the reader. Flan models seem to be better at processing the information from these non-gold passages than other models. As evidence, for NQ and k = 5, for the slice of data where there are no gold passages (paragraphs) but there are passages from the same Wikipedia pages as the gold passages, Flan models’ accuracy averages at 20%, which is much higher than LLaMa2 models’ average accuracy of 8%, LLaMa3 models’ at 4%, GPT-3.5’s at 10%, and Claude Haiku’s at 3%.
For BioASQ (Figure 7), one key difference is that GPT’s top-k performance is above its no-context performance for and LLaMa2 7B’s too for all , as opposed to being consistently under their no-context baseline. This indicates that for special-domain questions, these models may have stronger guardrails against misleading or irrelevant information, and can even fall back to using its pre-trained knowledge. Full results for other models are in Figure 12.
7 Impact of Retriever Choice
@@
We compare different retrievers and study their influence on the reader model’s downstream performance. Our RAGGED analysis guidelines are:
@@
Neural retrievers generally perform better than lexical retrievers. How much that advantage translates to a gain in reader performance highly depends on the domain and the complexity of the question (number of hops). We find that such advantages are retained for single-hop, open-domain questions, dampened for multi-hop, open-domain questions, and amplified for special-domain questions.
@@
These analyses inform which retrievers perform better and how much that translates to an improvement in reader performance. This can help practitioners decide which retriever is worth investing in for their specific tasks.
Open Domain, Single-Hop
For NQ questions, using ColBERT offers substantial improvements over using BM25 (19-point recall@k gain). Taking a look at the example at , the reader performance paired with ColBERT is strictly better than when paired with BM25, and that retriever gain translates to a reader gain of 9 F1 points when averaged over all models.
Open-Domain, Multi-Hop
Although ColBERT outperforms BM25 significantly (15-point recall@k gain), its impact on reader performance is much smaller. Taking a look at the example at , the average reader performance with CoLBERT only offers a 4-point F1 gain. This suggests that the challenge with multi-hop questions lies more with the reader’s multi-hop reasoning capability than with the retriever.
Biomedical Domain
ColBERT only outperforms BM25 minimally on BioASQ (<1 point), yet still offers a 4-point F1 gain in reader performance. This shows how a small retriever performance gain can be amplified in the reader performance under a special domain that the reader may have encountered less in pretraining.
@@ Appendix A Implementation Details
We truncate the Context to make sure the the rest of the prompt still fits within a reader’s context limit. Specifically, when using FlanT5 and FlanUL2 readers, we use T5Tokenizer to truncate sequences to up to 2 tokens; when using LLaMa models, we apply the LlamaTokenizer and truncate sequences by 4 tokens for LLaMa2 and 8 for LLaMa3. For closed-source models, we spent around $300. Subsequently, we incorporate a concise question-and-answer format that segments the query using "Question:" and cues the model’s response with "Answer:", ensuring precise and targeted answers.
For our reader decoding strategy, we used greedy decoding with a beam size of 1 and temperature of 1, selecting the most probable next word at each step without sampling. The output generation was configured to produce responses with 10 tokens. The experiments were conducted on NVIDIA A6000 GPUs, supported by an environment with 60GB RAM. The average response time was 1.1s per query when processing with a batch size of 50.
@@ Appendix B Dataset Details
For NQ and HotpotQA datasets in the open domain, we use the Wikipedia paragraphs corpus provided by the KILT benchmark (Petroni et al., 2021). For BioASQ, we use the PubMed Annual Baseline Repository for 2023 (of Medicine, 2023), where each passage is either a title or an abstract of PubMed papers. Dataset sizes are in Table 2.
@@
Appendix C Retriever Performance
| Retriever | Recall@k | |||||
| 1 | 2 | 5 | 10 | 20 | 50 | |
| NQ | ||||||
| BM25 | 2.7 | 4.4 | 8.0 | 11.5 | 16.3 | 22.8 |
| 10.3 | 16.3 | 27.8 | 36.8 | 47.7 | 53.2 | |
| ColBERT | 12.3 | 18.0 | 25.7 | 32.1 | 38.1 | 41.8 |
| 27.2 | 38.8 | 54.4 | 65.0 | 72.9 | 77.2 | |
| HotpotQA | ||||||
| BM25 | 19.1 | 25.9 | 34.6 | 41.1 | 46.8 | 54.2 |
| 23.3 | 31.2 | 42.7 | 52.1 | 59.1 | 62.8 | |
| ColBERT | 31.1 | 40.1 | 49.9 | 56.2 | 61.9 | 64.9 |
| 34.2 | 44.7 | 56.3 | 63.6 | 69.9 | 73.1 | |
| BioASQ | ||||||
| BM25 | 8.8 | 12.9 | 19.6 | 25.8 | 33.3 | 37.8 |
| 12.4 | 16.4 | 23.9 | 30.6 | 38.7 | 43.6 | |
| ColBERT | 8.8 | 13.5 | 20.7 | 27.1 | 34.3 | 38.6 |
| 14.2 | 18.2 | 25.6 | 32.2 | 39.8 | 44.2 | |
Appendix D Additional Reader Results